OpenStudy (anonymous):

Find the solution to -5 (x+1) >20

OpenStudy (anonymous):

The choices are a . x > 5 b. x < 5 c. x > -5 d. x < -5

OpenStudy (anonymous):

I'm so sorry... I wrote positive five and it is a negative five. So, it should be like this. -5(x+1)>20 -5(x+1)=20 -5x-5=20 -5x=25 x=-5 x<-5 (because we divided by a negative)

OpenStudy (anonymous):

Thank you so much !